Directions: Define the variables and write an equation to model the relationship in each table.
(This is one table with 2 coulumns)
Number of sales Total earnings
5 $2.00
10 $4.00
15 $6.00
20 $8.00
Can you help me with this?

Answer by Earlsdon(6103)

First, define the variables:
Let E be total earnings and S be number of sales. Since the total earnings(E) depends on the number of sales (S) then (E) will be the "dependent variable" and (S) will be the "independent variable"
Now find the relationship between the sales figure (S) and the total earnings (E).
One way to do this is the divide the total earnings figure by the corresponding number of sales figure and see if there is a constant quotient.
$2.00/ 5 = $0.40
$4.00/10 = $0.40
$6.00/15 = $0.40
$8.00/20 = $0.40
So you can see from this that if you multiply the number of sales (S) by $0.40 you will get the total earnings figure (E).
Now you can write the equation that relates these two quantities:
E = ($0.40)S Or, in words, Total earnings equals the number of sales times $0.40